# Learning Theory for Kernel Bilevel Optimization

### v1

@@ Appendix F Differentiability Results
Under Assumptions (A) and (B), the following properties hold for $L$:
@@
  • The map $(\omega,h)\mapsto\partial_{h}L(\omega,h)$ is differentiable. Moreover, for any $(\omega,h)\in\mathcal{U}\times\mathcal{H}$, its partial derivatives $\partial_{\omega,h}^{2}L(\omega,h)$ and $\partial_{\omega,h}^{2}L(\omega,h)$ at $(\omega,h)$ are Hilbert-Schmidt operators given by:
@@
Finite values. Fix $\omega\in\mathcal{U}$ and $h\in\mathcal{H}$. We will first show that $h$ is bounded on $\mathcal{X}$. By the reproducing property, we know that ${\left\lvert h(x)\right\rvert}\leq{\left\lVert h\right\rVert}_{\mathcal{H}}\sqrt{K(x,x)}$ for any $x\in\mathcal{X}$. Moreover, the kernel $K$ is bounded by a constant $\kappa$ thanks to Assumption (A). Consequently, ${\left\lvert h(x)\right\rvert}$ is upper-bounded by ${\left\lVert h\right\rVert}_{\mathcal{H}}\sqrt{\kappa}$ for any $x\in\mathcal{X}$.
Denote by $\mathcal{I}$ the compact interval defined as $\mathcal{I}=[-{\left\lVert h\right\rVert}_{\mathcal{H}}\sqrt{\kappa},{\left\lVert h\right\rVert}_{\mathcal{H}}\sqrt{\kappa}]$. By Assumption (B), the set $\mathcal{Y}$ is compact so that ×$\mathcal{I}\times\mathcal{Y}$ is also compact. Moreover, we know, by assumption on $\ell$, that $(\omega,v,y)\mapsto\ell(\omega,v,y)$ is continuous on $\mathcal{U}\times\mathbb{R}\times\mathcal{Y}$. Therefore, $(v,y)\mapsto\ell(\omega,v,y)$ must be bounded by some finite constant $C$ on the compact set ×$\mathcal{I}\times\mathcal{Y}$. This allows to deduce that $(x,y)\mapsto\ell(\omega,h(x),y)$ is bounded by $C$ for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$ and a fortiori $\mathbb{D}$-integrable, which shows that $L(\omega,h)$ is finite.
@@
$r_{j}(x,y)=\frac{\ell\left(\omega_{j},h_{j}(x),y\right)-\ell\left(\omega,h(x),y\right)-\left\langle\partial_{v}\ell\left(\omega,h(x),y\right)K(x,\cdot),h_{j}-h\right\rangle_{\mathcal{H}}-\langle\partial_{\omega}\ell(\omega,h(x),y),\omega_{j}-\omega\rangle}{\left\|(\omega_{j},h_{j})-(\omega,h)\right\|}.$ (25)
@@
We will first show that $\mathbb{E}_{\mathbb{D}}{\left[{\left\lvert r_{j}(x,y)\right\rvert}\right]}$ convergence to 00 by the dominated convergence theorem (Rudin, 1987, Theorem 1.34). By the reproducing property, note that $\ell(\omega,h(x),y)=\ell(\omega,\langle h,K(x,\cdot)\rangle_{\mathcal{H}},y)$. Hence, since $\ell$ is jointly differentiable in $(\omega,v)$ for any $y$, it follows that $(\omega,h)\mapsto\ell(\omega,h(x),y)$ is also differentiable for any $(x,y)$ by composition with the evaluation map $(\omega,h)\mapsto(\omega,\langle h,K(x,\cdot)\rangle_{\mathcal{H}}$ which is differentiable. Hence, the sequence $r_{j}(x,y)$ converges to 00 for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Moreover, by the mean-value theorem, there exists $0\leq c_{j}\leq 1$ such that:
@@
$r_{j}(x,y)=\frac{\left\langle{\left(\partial_{v}\ell\left(\bar{\omega}_{j},\bar{h}_{j}(x),y\right)-\partial_{v}\ell\left(\omega,h(x),y\right)\right)}K(x,\cdot),h_{j}-h\right\rangle_{\mathcal{H}}-\langle\partial_{\omega}\ell\left(\bar{\omega}_{j},\bar{h}_{j}(x),y\right)-\partial_{\omega}\ell\left(\omega,h(x),y\right),\omega_{j}-\omega\rangle}{\left\|(\omega_{j},h_{j})-(\omega,h)\right\|},$
@@
where $(\bar{\omega}_{j},\bar{h}_{j})\coloneqq(1-c_{j})(\omega,h)+c_{j}(\omega_{j},h_{j})$. We will show that $r_{j}(x,y)$ are bounded for $j$ large enough. We first construct a compact set that will contain all elements of the form $(\omega_{j},h_{j}(x),y)$ and $(\bar{\omega}_{j},\bar{h}_{j}(x),y)$ for all $j$ large enough. Since $\omega$ is an element in the open set $\mathcal{U}$, there exists a closed ball $\mathcal{B}(\omega,R)$ centered in $\omega$ and with some radius $R$ small enough so that $\mathcal{B}(\omega,R)$ is included in $\mathcal{U}$. For all $j$ large enough, $\omega_{j}$ and $\bar{\omega}_{j}$ belong to $\mathcal{B}(\omega,R)$ as these sequences converge to $\omega$. Moreover, $h_{j}$ and $\bar{h}_{j}$ are convergent sequences. Consequently, they must be bounded by some constant $B$. By the reproducing property, and recalling that the kernel $K$ is bounded by $\kappa$ by Assumption (A), it follows that $\max({\left\lvert h_{j}(x)\right\rvert},{\left\lvert\bar{h}_{j}(x)\right\rvert})\leq B\kappa$. Consider now the set $\mathcal{W}\coloneqq\mathcal{B}(\omega,R)\times\mathcal{B}_{1}(0,B\kappa)\times\mathcal{Y}$ which is a product of compact sets (recalling that $\mathcal{Y}$ is compact by Assumption (B)), where $\mathcal{B}_{1}(0,B\kappa)$ is the closed ball in $\mathbb{R}$ centered at 00 and of radius $B\kappa$. For $j$ large enough, we have established that $(\omega_{j},h_{j}(x),y)$ and $(\bar{\omega}_{j},\bar{h}_{j}(x),y)$ belong to $\mathcal{W}$ for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Since, by assumption on $\ell$, $\partial_{v}\ell(\omega,v,y)$ and $\partial_{\omega}\ell(\omega,v,y)$ are continuous, they must be bounded on the compact set $\mathcal{W}$ by some constant $C$. This allows to deduce from the expression of $r_{j}(x,y)$ above that $r_{j}(x,y)$ is bounded, and a fortiori dominated by an integrable function (a constant function). We then deduce that $\mathbb{E}_{\mathbb{\mathbb{D}}}{\left[{\left\lvert r_{j}(x,y)\right\rvert}\right]}$ converges to 00 by application of the dominated convergence theorem (Rudin, 1987, Theorem 1.34).
@@
Recalling Equation 25, $\mathbb{E}_{\mathbb{\mathbb{D}}}{\left[r_{j}(x,y)\right]}$ admits the following expression:
@@
$\mathbb{E}_{\mathbb{\mathbb{D}}}{\left[r_{j}(x,y)\right]}=\frac{L(\omega_{j},h_{j})-L(\omega,h)-\mathbb{E}_{\mathbb{D}}{\left[\left\langle\partial_{v}\ell\left(\omega,h(x),y\right)K(x,\cdot),h_{j}-h\right\rangle_{\mathcal{H}}\right]}-\langle\mathbb{E}_{\mathbb{D}}{\left[\partial_{\omega}\ell(\omega,h(x),y)\right]},\omega_{j}-\omega\rangle}{\left\|(\omega_{j},h_{j})-(\omega,h)\right\|}.$ (26)
@@
The convergence to 00 of the above expression precisely means that $L$ is differentiable at $(\omega,h)$ provided that the linear form $g\mapsto\mathbb{E}_{\mathbb{D}}{\left[\left\langle\partial_{v}\ell\left(\omega,h(x),y\right)K(x,\cdot),g\right\rangle_{\mathcal{H}}\right]}$ is bounded. To establish this fact, consider the RKHS-valued function $(x,y)\mapsto\partial_{v}\ell\left(\omega,h(x),y\right)K(x,\cdot)$. This function is Bochner-integrable in the sense that $\mathbb{E}_{\mathbb{D}}{\left[{\left\lVert\partial_{v}\ell\left(\omega,h(x),y\right)K(x,\cdot)\right\rVert}_{\mathcal{H}}\right]}$ is finite (Diestel & Uhl, 1977, Definition 1, Chapter 2). Indeed, we have the following:
@@
$\mathbb{E}_{\mathbb{D}}{\left[{\left\lVert\partial_{v}\ell\left(\omega,h(x),y\right)K(x,\cdot)\right\rVert}_{\mathcal{H}}\right]}\coloneqq\mathbb{E}_{\mathbb{D}}{\left[{\left\lvert\partial_{v}\ell\left(\omega,h(x),y\right)\right\rvert}\sqrt{K(x,x)}\right]}\leq\sqrt{\kappa}\mathbb{E}_{\mathbb{D}}{\left[{\left\lvert\partial_{v}\ell\left(\omega,h(x),y\right)\right\rvert}\right]}<+\infty,$
@@
where, for the inequality, we used that $(x,y)\mapsto\partial_{v}\ell(\omega,h(x),y)$ is bounded as shown previously. Consequently, $\mathbb{E}{\left[\partial_{v}\ell\left(\omega,h(x),y\right)K(x,\cdot)\right]}$ is an element in $\mathcal{H}$ satisfying:
@@
$\langle\mathbb{E}{\left[\partial_{v}\ell\left(\omega,h(x),y\right)K(x,\cdot)\right]},g\rangle_{\mathcal{H}}=\mathbb{E}_{\mathbb{D}}{\left[\left\langle\partial_{v}\ell\left(\omega,h(x),y\right)K(x,\cdot),g\right\rangle_{\mathcal{H}}\right]},\forall g\in\mathcal{H}.$
@@
The above property follows from (Diestel & Uhl, 1977, Theorem 6, Chapter 2) for Bochner-integrable functions that allows exchanging the integral and the application of a continuous linear map (here the scalar product with an element $g$). The above identity establishes that $g\mapsto\mathbb{E}_{\mathbb{D}}{\left[\left\langle\partial_{v}\ell\left(\omega,h(x),y\right)K(x,\cdot),g\right\rangle_{\mathcal{H}}\right]}$ is bounded and provides the desired expression for $\partial_{h}L(\omega,h)$. The expression for $\partial_{\omega}L(\omega,h)$ directly follows from the last term in Equation 26.
@@
$\begin{split}\left\|(\omega_{j},h_{j})-(\omega,h)\right\|s_{j}(x,y)=&{\left(\partial_{v}\ell\left(\omega_{j},h_{j}(x),y\right)-\partial_{v}\ell\left(\omega,h(x),y\right)\right)}K(x,\cdot)\\ &-\partial_{v}^{2}\ell\left(\omega,h(x),y\right)K(x,\cdot)\otimes K(x,\cdot){\left(h_{j}-h\right)}-{\left(\omega_{j}-\omega\right)}^{\top}\partial_{\omega,v}^{2}\ell(\omega,h(x),y)K(x,\cdot).\end{split}$ (27)
We will first show that $\mathbb{E}_{\mathbb{D}}{\left[{\left\lVert s_{j}(x,y)\right\rVert}_{\mathcal{H}}\right]}$ convergence to 00 by the dominated convergence theorem for Bochner-integrable functions (Diestel & Uhl, 1977, Theorem 3, Chapter 2). By the reproducing property, note that $\partial_{v}\ell(\omega,h(x),y)K(x,\cdot)=\partial_{v}\ell(\omega,\langle h,K(x,\cdot)\rangle_{\mathcal{H}},y)K(x,\cdot)$. Hence, since $(\omega,v)\mapsto\partial_{v}\ell(\omega,v,y)$ is jointly differentiable in $(\omega,v)$ for any $y$, it follows that $(\omega,h)\mapsto\partial_{v}\ell(\omega,h(x),y)K(x,\cdot)$ is also differentiable for any $(x,y)$ by composition with the evaluation map $(\omega,h)\mapsto(\omega,\langle h,K(x,\cdot)\rangle_{\mathcal{H}}$ which is differentiable. Hence, the sequence $s_{j}(x,y)$ converges to 00 for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Moreover, by the mean-value theorem, there exists $0\leq c_{j}\leq 1$ such that:
$\begin{split}\left\|(\omega_{j},h_{j})-(\omega,h)\right\|s_{j}(x,y)=&\partial_{v}^{2}\ell\left(\bar{\omega}_{j},\bar{h}_{j}(x),y\right)K(x,\cdot)\otimes K(x,\cdot){\left(h_{j}-h\right)}+{\left(\omega_{j}-\omega\right)}^{\top}\partial_{\omega,v}^{2}\ell(\bar{\omega}_{j},\bar{h}_{j}(x),y)K(x,\cdot)\\ &-\partial_{v}^{2}\ell\left(\omega,h(x),y\right)K(x,\cdot)\otimes K(x,\cdot){\left(h_{j}-h\right)}-{\left(\omega_{j}-\omega\right)}^{\top}\partial_{\omega,v}^{2}\ell(\omega,h(x),y)K(x,\cdot),\end{split}$
@@
This already establishes that $s_{j}(x,y)$ is bounded so that $\mathbb{E}_{\mathbb{D}}{\left[{\left\lVert s_{j}(x,y)\right\rVert}_{\mathcal{H}}\right]}$ converges to 00 by application of the dominated convergence theorem. Recalling Equation 27, $\mathbb{E}_{\mathbb{\mathbb{D}}}{\left[s_{j}(x,y)\right]}$ admits the following expression:
$\begin{split}\left\|(\omega_{j},h_{j})-(\omega,h)\right\|\mathbb{E}_{\mathbb{\mathbb{D}}}{\left[s_{j}(x,y)\right]}=&\partial_{h}L(\omega_{j},h_{j})-\partial_{h}L(\omega,h)-\mathbb{E}_{\mathbb{\mathbb{D}}}{\left[\partial_{v}^{2}\ell\left(\omega,h(x),y\right)K(x,\cdot)\otimes K(x,\cdot){\left(h_{j}-h\right)}\right]}\\ &-{\left(\omega_{j}-\omega\right)}^{\top}\mathbb{E}_{\mathbb{\mathbb{D}}}{\left[\partial_{\omega,v}^{2}\ell(\omega,h(x),y)K(x,\cdot)\right]}.\end{split}$
The convergence to 00 of the above expression precisely means that $L$ is differentiable at $(\omega,h)$ provided that: (1) $\mathbb{E}_{\mathbb{\mathbb{D}}}{\left[\partial_{\omega,v}^{2}\ell(\omega,h(x),y)K(x,\cdot)\right]}$ is an element in $\mathcal{H}^{d}$, and (2) the linear map $g\mapsto\mathbb{E}_{\mathbb{D}}{\left[\partial_{v}^{2}\ell\left(\omega,h(x),y\right)(K(x,\cdot)\otimes K(x,\cdot))g\right]}$ is bounded. Using the same strategy to establish Bochner’s inetgrability of $(x,y)\mapsto\partial_{v}\ell\left(\omega,h(x),y\right)K(x,\cdot)$, we can show that $(x,y)\mapsto\partial_{\omega,v}^{2}\ell\left(\omega,h(x),y\right)K(x,\cdot)$ is also Bochner-integrable so that $\mathbb{E}_{\mathbb{\mathbb{D}}}{\left[\partial_{\omega,v}^{2}\ell(\omega,h(x),y)K(x,\cdot)\right]}$ is indeed an element in $\mathcal{H}^{d}$. This also establishes the expression of $\partial_{\omega,h}L(\omega,h)$. Similarly, we consider the operator-valued function $\xi:(x,y)\mapsto\partial_{v}^{2}\ell\left(\omega,h(x),y\right)K(x,\cdot)\otimes K(x,\cdot)$ with values in the space of Hilbert-Schmidt operators on $\mathcal{H}$. The Hilbert-Schmidt (HS) norm of such function satisfies the following inequality:
@@
The above property follows from (Diestel & Uhl, 1977, Theorem 6, Chapter 2) for Bochner-integrable functions that allows exchanging the integral and the application of a continuous linear map (here the scalar product with an element $g$). Hence, from the above identity we deduce the desired expression for $\partial_{h}^{2}L(\omega,h)$. ∎
Appendix G Auxiliary Technical Lemmas
Lemma G.1.

### v3

@@ Appendix G Differentiability Results
Under Assumptions˜(A), (B) and (C), the following properties hold for $L$:
@@
  • The map $(\omega,h)\mapsto\partial_{h}L(\omega,h)$ is differentiable. Moreover, for any $(\omega,h)\in\mathcal{U}\times\mathcal{H}$, its partial derivatives $\partial_{\omega,h}^{2}L(\omega,h)$ and $\partial_{h}^{2}L(\omega,h)$ at $(\omega,h)$ are Hilbert-Schmidt operators given by:
@@
Finite values. Fix $\omega\in\mathcal{U}$ and $h\in\mathcal{H}$. We will first show that $h$ is bounded on $\mathcal{X}$. By the reproducing property, we know that $\left\lvert h(x)\right\rvert\leq\left\lVert h\right\rVert_{\mathcal{H}}\sqrt{K(x,x)}$ for any $x\in\mathcal{X}$. Moreover, the kernel $K$ is bounded by a constant $\kappa$ thanks to Assumption˜(B). Consequently, $\left\lvert h(x)\right\rvert$ is upper-bounded by $\left\lVert h\right\rVert_{\mathcal{H}}\sqrt{\kappa}$ for any $x\in\mathcal{X}$.
Denote by $\mathcal{I}$ the compact interval defined as $\mathcal{I}=[-\left\lVert h\right\rVert_{\mathcal{H}}\sqrt{\kappa},\left\lVert h\right\rVert_{\mathcal{H}}\sqrt{\kappa}]$. By Assumption˜(C), the set $\mathcal{Y}$ is compact so that $\mathcal{I}\times\mathcal{Y}$ is also compact. Moreover, we know, by assumption on $\ell$, that $(\omega,v,y)\mapsto\ell(\omega,v,y)$ is continuous on $\mathcal{U}\times\mathbb{R}\times\mathcal{Y}$. Therefore, $(v,y)\mapsto\ell(\omega,v,y)$ must be bounded by some finite constant $C$ on the compact set $\mathcal{I}\times\mathcal{Y}$. This allows to deduce that $(x,y)\mapsto\ell(\omega,h(x),y)$ is bounded by $C$ for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$ and a fortiori $\mathbb{D}$-integrable, which shows that $L(\omega,h)$ is finite.
@@
$\frac{\ell\left(\omega_{j},h_{j}(x),y\right)-\ell\left(\omega,h(x),y\right)-\left\langle\partial_{v}\ell\left(\omega,h(x),y\right)K(x,\cdot),h_{j}-h\right\rangle_{\mathcal{H}}-\langle\partial_{\omega}\ell(\omega,h(x),y),\omega_{j}-\omega\rangle}{\left\|(\omega_{j},h_{j})-(\omega,h)\right\|}.$ (24)
@@
We will first show that $\mathbb{E}_{\mathbb{D}}\left[\left\lvert r_{j}(x,y)\right\rvert\right$] converges to 0 by the dominated convergence theorem [rudin1987real, Theorem 1.34]. By the reproducing property, note that $\ell(\omega,h(x),y)=\ell(\omega,\langle h,K(x,\cdot)\rangle_{\mathcal{H}},y)$. Hence, since $\ell$ is jointly differentiable in $(\omega,v)$ for any $y$, it follows that $(\omega,h)\mapsto\ell(\omega,h(x),y)$ is also differentiable for any $(x,y)$ by composition with the evaluation map $(\omega,h)\mapsto(\omega,\langle h,K(x,\cdot)\rangle_{\mathcal{H}}$ which is differentiable. Hence, the sequence $r_{j}(x,y)$ converges to 0 for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Moreover, by the mean-value theorem, there exists $0\leq c_{j}\leq 1$ such that:
$r_{j}(x,y)=$
@@
$\frac{\left\langle\left(\partial_{v}\ell\left(\bar{\omega}_{j},\bar{h}_{j}(x),y\right)-\partial_{v}\ell\left(\omega,h(x),y\right)\right)K(x,\cdot),h_{j}-h\right\rangle_{\mathcal{H}}-\langle\partial_{\omega}\ell\left(\bar{\omega}_{j},\bar{h}_{j}(x),y\right)-\partial_{\omega}\ell\left(\omega,h(x),y\right),\omega_{j}-\omega\rangle}{\left\|(\omega_{j},h_{j})-(\omega,h)\right\|},$
@@
where $(\bar{\omega}_{j},\bar{h}_{j})\coloneqq(1-c_{j})(\omega,h)+c_{j}(\omega_{j},h_{j})$. We will show that $r_{j}(x,y)$ is bounded for $j$ large enough. We first construct a compact set that will contain all elements of the form $(\omega_{j},h_{j}(x),y)$ and $(\bar{\omega}_{j},\bar{h}_{j}(x),y)$ for all $j$ large enough. Since $\omega$ is an element in the open set $\mathcal{U}$, there exists a closed ball $\mathcal{B}(\omega,R)$ centered at $\omega$ and with some radius $R$ small enough so that $\mathcal{B}(\omega,R)$ is included in $\mathcal{U}$. For all $j$ large enough, $\omega_{j}$ and $\bar{\omega}_{j}$ belong to $\mathcal{B}(\omega,R)$ as these sequences converge to $\omega$. Moreover, $h_{j}$ and $\bar{h}_{j}$ are convergent sequences. Consequently, they must be bounded by some constant $B$. By the reproducing property, and recalling that the kernel $K$ is bounded by $\kappa$ by Assumption˜(B), it follows that $\max(\left\lvert h_{j}(x)\right\rvert,\left\lvert\bar{h}_{j}(x)\right\rvert)\leq B\kappa$. Consider now the set $\mathcal{W}\coloneqq\mathcal{B}(\omega,R)\times\mathcal{B}_{1}(0,B\kappa)\times\mathcal{Y}$ which is a product of compact sets (recalling that $\mathcal{Y}$ is compact by Assumption˜(C)), where $\mathcal{B}_{1}(0,B\kappa)$ is the closed ball in $\mathbb{R}$ centered at 0 and of radius $B\kappa$. For $j$ large enough, we have established that $(\omega_{j},h_{j}(x),y)$ and $(\bar{\omega}_{j},\bar{h}_{j}(x),y)$ belong to $\mathcal{W}$ for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Since, by assumption on $\ell$, $\partial_{v}\ell(\omega,v,y)$ and $\partial_{\omega}\ell(\omega,v,y)$ are continuous, they must be bounded on the compact set $\mathcal{W}$ by some constant $C$. This allows to deduce from the expression of $r_{j}(x,y)$ above that $r_{j}(x,y)$ is bounded, and a fortiori dominated by an integrable function (a constant function). We then deduce that $\mathbb{E}_{\mathbb{\mathbb{D}}}\left[\left\lvert r_{j}(x,y)\right\rvert\right$] converges to 0 by application of the dominated convergence theorem [rudin1987real, Theorem 1.34].
@@
Recalling Appendix˜G, $\mathbb{E}_{\mathbb{\mathbb{D}}}\left[r_{j}(x,y)\right]$ admits the following expression:
$\mathbb{E}_{\mathbb{\mathbb{D}}}\left[r_{j}(x,y)\right]=$
@@
$\frac{L(\omega_{j},h_{j})-L(\omega,h)-\mathbb{E}_{\mathbb{D}}\left[\left\langle\partial_{v}\ell\left(\omega,h(x),y\right)K(x,\cdot),h_{j}-h\right\rangle_{\mathcal{H}}\right]-\langle\mathbb{E}_{\mathbb{D}}\left[\partial_{\omega}\ell(\omega,h(x),y)\right],\omega_{j}-\omega\rangle}{\left\|(\omega_{j},h_{j})-(\omega,h)\right\|}.$ (25)
@@
The convergence to 0 of the above expression precisely means that $L$ is differentiable at $(\omega,h)$ provided that the linear form $g\mapsto\mathbb{E}_{\mathbb{D}}\left[\left\langle\partial_{v}\ell\left(\omega,h(x),y\right)K(x,\cdot),g\right\rangle_{\mathcal{H}}\right]$ is bounded. To establish this fact, consider the RKHS-valued function $(x,y)\mapsto\partial_{v}\ell\left(\omega,h(x),y\right)K(x,\cdot)$. This function is Bochner-integrable in the sense that $\mathbb{E}_{\mathbb{D}}\left[\left\lVert\partial_{v}\ell\left(\omega,h(x),y\right)K(x,\cdot)\right\rVert_{\mathcal{H}}\right]$ is finite [diestel1977vector, Definition 1, Chapter 2]. Indeed, we have the following:
@@
$\mathbb{E}_{\mathbb{D}}\left[\left\lVert\partial_{v}\ell\left(\omega,h(x),y\right)K(x,\cdot)\right\rVert_{\mathcal{H}}\right]$ $\coloneqq\mathbb{E}_{\mathbb{D}}\left[\left\lvert\partial_{v}\ell\left(\omega,h(x),y\right)\right\rvert\sqrt{K(x,x)}\right]$
$\leq\sqrt{\kappa}\mathbb{E}_{\mathbb{D}}\left[\left\lvert\partial_{v}\ell\left(\omega,h(x),y\right)\right\rvert\right$$\infty$,
@@
where, for the inequality, we used that $(x,y)\mapsto\partial_{v}\ell(\omega,h(x),y)$ is bounded as shown previously. Consequently, $\mathbb{E}_{\mathbb{D}}\left[\partial_{v}\ell\left(\omega,h(x),y\right)K(x,\cdot)\right]$ is an element in $\mathcal{H}$ satisfying:
@@
$\langle\mathbb{E}_{\mathbb{D}}\left[\partial_{v}\ell\left(\omega,h(x),y\right)K(x,\cdot)\right],g\rangle_{\mathcal{H}}=\mathbb{E}_{\mathbb{D}}\left[\left\langle\partial_{v}\ell\left(\omega,h(x),y\right)K(x,\cdot),g\right\rangle_{\mathcal{H}}\right],\forall g\in\mathcal{H}.$
@@
The above property follows from [diestel1977vector, Theorem 6, Chapter 2] for Bochner-integrable functions that allows exchanging the integral and the application of a continuous linear map (here the scalar product with an element $g$). The above identity establishes that $g\mapsto\mathbb{E}_{\mathbb{D}}\left[\left\langle\partial_{v}\ell\left(\omega,h(x),y\right)K(x,\cdot),g\right\rangle_{\mathcal{H}}\right]$ is bounded and provides the desired expression for $\partial_{h}L(\omega,h)$. The expression for $\partial_{\omega}L(\omega,h)$ directly follows from the last term in Appendix˜G.
@@
$\begin{split}\left\|(\omega_{j},h_{j})-(\omega,h)\right\|s_{j}(x,y)=&\left(\partial_{v}\ell\left(\omega_{j},h_{j}(x),y\right)-\partial_{v}\ell\left(\omega,h(x),y\right)\right)K(x,\cdot)\\ &-\partial_{v}^{2}\ell\left(\omega,h(x),y\right)K(x,\cdot)\otimes K(x,\cdot)\left(h_{j}-h\right)\\ &-\left(\omega_{j}-\omega\right)^{\top}\partial_{\omega,v}^{2}\ell(\omega,h(x),y)K(x,\cdot).\end{split}$ (26)
We will first show that $\mathbb{E}_{\mathbb{D}}\left[\left\lVert s_{j}(x,y)\right\rVert_{\mathcal{H}}\right]$ converges to 0 by the dominated convergence theorem for Bochner-integrable functions [diestel1977vector, Theorem 3, Chapter 2]. By the reproducing property, note that $\partial_{v}\ell(\omega,h(x),y)K(x,\cdot)=\partial_{v}\ell(\omega,\langle h,K(x,\cdot)\rangle_{\mathcal{H}},y)K(x,\cdot)$. Hence, since $(\omega,v)\mapsto\partial_{v}\ell(\omega,v,y)$ is jointly differentiable in $(\omega,v)$ for any $y$, it follows that $(\omega,h)\mapsto\partial_{v}\ell(\omega,h(x),y)K(x,\cdot)$ is also differentiable for any $(x,y)$ by composition with the evaluation map $(\omega,h)\mapsto(\omega,\langle h,K(x,\cdot)\rangle_{\mathcal{H}}$ which is differentiable. Hence, the sequence $s_{j}(x,y)$ converges to 0 for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Moreover, by the mean-value theorem, there exists $0\leq c_{j}\leq 1$ such that:
$\begin{split}\left\|(\omega_{j},h_{j})-(\omega,h)\right\|s_{j}(x,y)=&\partial_{v}^{2}\ell\left(\bar{\omega}_{j},\bar{h}_{j}(x),y\right)K(x,\cdot)\otimes K(x,\cdot)\left(h_{j}-h\right)\\ &+\left(\omega_{j}-\omega\right)^{\top}\partial_{\omega,v}^{2}\ell(\bar{\omega}_{j},\bar{h}_{j}(x),y)K(x,\cdot)\\ &-\partial_{v}^{2}\ell\left(\omega,h(x),y\right)K(x,\cdot)\otimes K(x,\cdot)\left(h_{j}-h\right)\\ &-\left(\omega_{j}-\omega\right)^{\top}\partial_{\omega,v}^{2}\ell(\omega,h(x),y)K(x,\cdot),\end{split}$
@@
This already establishes that $s_{j}(x,y)$ is bounded so that $\mathbb{E}_{\mathbb{D}}\left[\left\lVert s_{j}(x,y)\right\rVert_{\mathcal{H}}\right]$ converges to 0 by application of the dominated convergence theorem. Recalling Equation˜26, $\mathbb{E}_{\mathbb{\mathbb{D}}}\left[s_{j}(x,y)\right]$ admits the following expression:
$\begin{split}\left\|(\omega_{j},h_{j})-(\omega,h)\right\|\mathbb{E}_{\mathbb{\mathbb{D}}}\left[s_{j}(x,y)\right]=&\partial_{h}L(\omega_{j},h_{j})-\partial_{h}L(\omega,h)\\ &-\mathbb{E}_{\mathbb{\mathbb{D}}}\left[\partial_{v}^{2}\ell\left(\omega,h(x),y\right)K(x,\cdot)\otimes K(x,\cdot)\left(h_{j}-h\right)\right]\\ &-\left(\omega_{j}-\omega\right)^{\top}\mathbb{E}_{\mathbb{\mathbb{D}}}\left[\partial_{\omega,v}^{2}\ell(\omega,h(x),y)K(x,\cdot)\right].\end{split}$
The convergence to 0 of the above expression precisely means that $L$ is differentiable at $(\omega,h)$ provided that: (1) $\mathbb{E}_{\mathbb{\mathbb{D}}}\left[\partial_{\omega,v}^{2}\ell(\omega,h(x),y)K(x,\cdot)\right]$ is an element in $\mathcal{H}^{d}$, and (2) the linear map $g\mapsto\mathbb{E}_{\mathbb{D}}\left[\partial_{v}^{2}\ell\left(\omega,h(x),y\right)(K(x,\cdot)\otimes K(x,\cdot))g\right]$ is bounded. Using the same strategy to establish Bochner’s integrability of $(x,y)\mapsto\partial_{v}\ell\left(\omega,h(x),y\right)K(x,\cdot)$, we can show that $(x,y)\mapsto\partial_{\omega,v}^{2}\ell\left(\omega,h(x),y\right)K(x,\cdot)$ is also Bochner-integrable so that $\mathbb{E}_{\mathbb{\mathbb{D}}}\left[\partial_{\omega,v}^{2}\ell(\omega,h(x),y)K(x,\cdot)\right]$ is indeed an element in $\mathcal{H}^{d}$. This also establishes the expression of $\partial_{\omega,h}L(\omega,h)$. Similarly, we consider the operator-valued function $\xi:(x,y)\mapsto\partial_{v}^{2}\ell\left(\omega,h(x),y\right)K(x,\cdot)\otimes K(x,\cdot)$ with values in the space of Hilbert-Schmidt operators on $\mathcal{H}$. The Hilbert-Schmidt (HS) norm of such function satisfies the following inequality:
@@
The above property follows from [diestel1977vector, Theorem 6, Chapter 2] for Bochner-integrable functions that allows exchanging the integral and the application of a continuous linear map (here the scalar product with an element $g$). Hence, from the above identity, we deduce the desired expression for $\partial_{h}^{2}L(\omega,h)$. ∎
Appendix H Auxiliary Technical Lemmas
Lemma H.1.